An ideal gas at 0.80 atmospheres and 87oC occupies 0.450 liter. How many moles are in the sample? (R =0.0821 liter-atm/mole-K)


1
Expert's answer
2021-10-22T03:19:22-0400

According to ideal gas law "p V = \nu R T", from where the amount of substance is "\nu = \frac{p V}{R T} = \frac{0.8 atm \cdot 0.45 l}{0.0821 \frac{l \cdot atm}{mol \cdot K} \cdot 360.15 K} \approx 0.012 mol".